OpenStudy (anonymous):

I need help solving 47

OpenStudy (kkutie7):

of?

OpenStudy (anonymous):

The problem 47 in the attachment

OpenStudy (anonymous):

@Kkutie7

OpenStudy (kkutie7):

there is no attachment

OpenStudy (anonymous):

Hhckvdhvg

OpenStudy (anonymous):

Now @Kkutie7

OpenStudy (kkutie7):

ok f(3)=2 and g(3)=1 2+1= 3

geerky42 (geerky42):

\[(f+g)(3) = f(3)+g(3)\]So you know that \(f(3)=2\) and \(g(3)=1\) So you have \(f+g)(3) = f(3)+g(3)=2+1=\boxed?\)